Rationalize the denominator:1. 5+3√2/5-3√2
1
Expert's answer
2012-06-12T08:33:09-0400
(5+3√2)/(5-3√2)&
Let's multiply numerator and denominator by 5+3√2:
(5+3√2)^2/{(5-3√2)(5+3√2)}=(5+3√2)^2/(25-18)=(25+30√2+18)/7=(43+30√2)/7